### v1

# The Remarkable Robustness of LLMs: Stages of Inference?

Abstract
We demonstrate and investigate the remarkable robustness of Large Language Models by deleting and swapping adjacent layers. We find that deleting and swapping interventions retain 72-95% of the original model’s prediction accuracy without fine-tuning, whereas models with more layers exhibit more robustness. Based on the results of the layer-wise intervention and further experiments, we hypothesize the existence of four universal stages of inference across eight different models: detokenization, feature engineering, prediction ensembling, and residual sharpening. The first stage integrates local information, lifting raw token representations into higher-level contextual representations. Next is the iterative refinement of task and entity-specific features. Then, the second half of the model begins with a phase transition, where hidden representations align more with the vocabulary space due to specialized model components. Finally, the last layer sharpens the following token distribution by eliminating obsolete features that add noise to the prediction.
1 Introduction
Recent advancements in Large Language Models (LLMs) have demonstrated remarkable reasoning capabilities, often attributed to their increased scale [67]. However, the benefits of scaling are accompanied by heightened risks and vulnerabilities [7, 36, 4], necessitating extensive research into the underlying mechanisms of these capabilities. Inspired by previous studies on model robustness [28, 45, 58, 44, 5, 65], this work investigates the sensitivity of LLMs to the deletion and swapping of entire layers during inference. Our findings suggest four universal stages of inference: detokenization, feature engineering, prediction ensembling, and residual sharpening.
Recent work in mechanistic interpretability has explored the iterative inference hypothesis [3, 58], which suggests that each layer incrementally updates the hidden state of a token in a direction of decreasing loss by gradually shaping the next token distribution [24]. Self-repair [58] and redundancy [45, 28] in networks further support this hypothesis of iterative inference. However, recent work also indicates a degree of specialization in networks, with attention heads and neurons playing specific roles [32, 43, 26], which compose into more sophisticated circuits [52, 21].
In this work, we begin by exploring the robustness of language models by performing a series of interventions that delete individual layers or swap adjacent layers (Figure 2). Using these results, we then attempt to understand the roles of different depths in the network. Our experiments suggest four phases in a model, which we investigate further.
Specifically, we hypothesize an initial (1) detokenization [15] stage, where the model integrates local context to convert raw token representations into coherent entities, as suggested by the sensitivity to deletion and swapping. In the (2) feature engineering stage, the model iteratively builds feature representations based on token context, leading to minimal progress in token prediction but significant increases in probing accuracy and patching importance. A phase transition from attention-heavy computation to MLP-heavy computation delineates the following stage. During the (3) prediction ensembling stage, the model emphasizes relevant predictions while suppressing others, potentially marked by high MLP computation and the emergence of prediction neurons. Finally, in the (4) residual sharpening stage, tokens transition from semantic representations to specific next-token predictions, again showing sensitivity to deletion and swapping. This framework represents a tentative step toward understanding the complex nature of token processing in advanced language models (see Table 1).
Figure 1 delineates our four characteristic phases, which we describe further in Table 1. We integrate our work with previous depth-related findings within interpretability [21].
| Stage | Name | Function | Observable signatures |
| 1 | Detokenization | Integrate local context to transform raw token representations into coherent entities | Catastrophic sensitivity to deletion and swapping |
| 2 | Feature Engineering | Iteratively build feature representation depending on token context | Little progress made towards next token prediction, but significant increase in probing accuracy and patching importance. Attention Heavy Computation |
| 3 | Prediction Ensembling | Convert previously constructed semantic features into plausible next token predictions using an ensemble of model components. | Increased MLP importance; prediction neurons appear; phase transition in progress towards final prediction |
| 4 | Residual Sharpening | Sharpen the next token distribution by eliminating obsolete features that add noise to the prediction | More suppression neurons than prediction neurons |
2 Related Work
Universal Mechanisms
A key activity of mechanistic interpretability is circuit analysis, where research uncovers relevant model components for a given computation. In computer vision, circuits discover how features are constructed across many layers [51]. Follow-up work found that feature building was carried out by specific mechanisms that appeared across models, such as frequency detectors [60] and curve-circuits [8]. Language models seem to be following a similar line of inquiry, first uncovering universal model components, such as induction heads [52], successor heads [26], and copy suppression [43] in attention mechanisms. The discovery of knowledge neurons [10] paved the way for the identification of various specialized neurons [32, 66]. These specialized components can be connected to critical roles in universal processes in language models, such as circuit reuse [47], variable finding mechanisms [19], self-repair [58, 44] (which also studies layer-wise ablations), function vectors [63, 35], and long context retrieval [64].
Depth-Dependent Studies
Circuit analysis naturally motivated various depth-dependent studies, such as the logit lens [50], a technique that reveals the model’s prediction distribution at each layer. This line of reasoning revived the iterative inference hypothesis, the idea that each layer updates the hidden state in a direction of decreasing loss, in the context of ResNets [27, 37]. Due to residual connections, researchers were able to demonstrate that models exhibited “ensembling" [65] through layer ablations and permutations, referred to as “lesion" studies. This was later applied to modern transformers [13, 5]. Recent work [3] expanded on the logit lens and provided further evidence for iterative inference. This hypothesis was also supported by analyzing transformers in embedding spaces [12] and model self-repair [58].
Many works have localized specific kinds of computations to particular regions of large language models. For instance, model editing research [46] showed that knowledge is stored in mid-layer MLP neurons. Follow-up studies [33, 25] suggested that these facts can be stored across layers, with different components encoding task-specific instructions and recall. There is significant literature citing changes halfway through a language model, such as linear probe quality improving the fastest in the first half of a model [30] or fine-tuning predominantly updating weights in the middle of a model [53]. Few works observed that activation sparsity changes in the middle of a model, transitioning from sparse to dense [40, 66].
Other works have extended these insights into phases, such as identifying emergent specific phases of truth processing in language models [41] and distilling translation in multilingual transformers into three distinct phases: input space, concept space, and output space [68]. Other works hypothesized stages of inference in large language models [15]. This was further studied in the context of iterative inference [3], which demonstrated the importance of the first layers through layer ablations. When permuting the layers of GPT-2 style transformers, studies found the best performance in variants that have a higher proportion of self-attention layers at the beginning of models and more feedforward layers at the end [54]. These studies thematically suggest a preferential order to computation, which we investigate further.
Robustness: Quantization, Pruning, Redundancy
Numerous studies have inadvertently revealed depth-related findings while exploring model pruning or quantization. Prior work on zero-ablations of transformers has predominantly concentrated on BERT [14] style transformers [71, 17, 18, 59, 69]. Despite significant differences in these models [16], many underlying principles may be applicable. For example, pruning the final layers of a model retains most of the model’s performance [59]. Follow-up studies suggest 85% neuron redundancy in BERT transformers [11]. This is analogous to recent findings in modern networks, which demonstrate that approximately 70% of attention heads and 20% of feed-forward networks can be removed with minimal impact on task performance [1], suggesting redundancy [28, 45]. Model quantization work found improved benchmark performance by only keeping low-rank components of MLPs in the second half of the models [61] while uncovering token frequency dependencies in MLP weights. While we only touch upon the relationship between robustness and token frequency, [42] proposes that pretraining data contributes to this robustness.
3 Experimental Protocol
Models
To investigate the stages of inference in language models, we examine the Pythia [6], GPT-2 [56], and Microsoft Phi [29, 39] model families, which range from 124M to 6.9B parameters (see Table 2). All families use decoder-only transformers but exhibit several architectural differences that enable us to test the generality of our findings. Pythia models utilize parallel attention and MLPs, executing these components concurrently during inference. In contrast, GPT-2 and Phi models apply these components sequentially, with attention followed by the MLP (see Figure 2). Additionally, GPT-2 models were trained with dropout. We preprocess weights identically across all models, folding in the layer norm, centering the unembedding weights, and centering the writing weight. Despite the architectural differences, these models exhibit consistent inference patterns that validate our hypothesis.
Data
We evaluate all three model families on a corpus of one million tokens from random sequences of the Pile dataset [22]. The Pile was used to train the Pythia models and includes OpenWebText, the training corpus for GPT-2. Testing models on data similar to their training data ensures a fair comparison and minimizes the impact of domain shifts on observed inference patterns.
| Parameters | Layers |
| Pythia (410M) | 24 |
| Pythia (1.4B) | 24 |
| Pythia (2.8B) | 32 |
| Pythia (6.9B) | 32 |
| Parameters | Layers |
| Small (124M) | 12 |
| Medium (355M) | 24 |
| Large (774M) | 36 |
| XL (1.5B) | 48 |
| Parameters | Layers |
| Phi 1 (1.3B) | 24 |
| Phi 1.5 (1.3B) | 24 |
| Phi 2 (2.7B) | 32 |
Layer Swap Data Collection
To study the robustness and role of different model components at different depths, we employ a swapping intervention where we switch the execution order of a pair of adjacent layers in the model. Specifically, for a swap intervention at layer , we execute the transformer block (including the attention layer, MLP, and normalization) before executing block . We record the Kullback-Leibler (KL) divergence between the intervened and original models, measuring the difference in their output distributions, along with model-wise metrics such as loss, top-1 prediction accuracy, and prediction entropy. This intervention allows us to examine how the order of computation affects the model’s behavior and performance at different depths.
Ablation Data Collection
To generate baselines for each layer swap experiment, we perform zero ablations on the corresponding layer while collecting the same metrics. The ablation preserves the swap ordering: for a swap ordering of 1-2-4-3-5, the ablation maintains 1-2-4-5. We opt for zero ablation as opposed to mean ablation, as proposed by [3], to maintain consistency with the swap order. Additionally, we perform attention-only and MLP-only ablations to study the specific roles of these components in the model’s inference process. By comparing the effects of layer swapping and ablation, we can gain insights into the importance and function of each component at different depths in the model.
4 Robustness
4.1 Intervention Results
To study the robustness of language models, we apply our aforementioned drop and swap interventions to every layer of four GPT-2 models [55] and four Pythia [6]. In Figure 3, we report (1) the KL divergence between the prediction of the intervened model and the nominal model, (2) the fraction of predictions that are the same between the intervened model and the baseline model (denoted as relative accuracy), and (3) the change in entropy of the prediction between the intervened and baseline model for all interventions.
Our results show that intervening on the first layer is catastrophic for model performance. Specifically, dropping or swapping the first layer causes the model to have very high entropy predictions as opposed to causing a mode collapse on a constant token. In Pythia models, swapping the last layer with the second to last layer also has a similar catastrophic high-entropy effect, while GPT-2 models largely preserve their predictions. We further discuss our hypothesized role of the first and last layer in Section 5.1 and Section 5.4 respectively.
In contrast to the first and last layer interventions, the middle layers are remarkably robust to both deletion and minor order changes. When zooming in on the differences between the effect of swaps and drops for intermediate layers, we find that swapping adjacent layers is less harmful than ablating layers. These results match similar experiments performed on vision transformers [5]. We take this as evidence that certain operations within the forward pass are commutative, though further experimentation is required.
We suspect that GPT-2 exhibits greater robustness than Pythia because (1) GPT-2 models are trained with dropout, which likely increases redundancy and (2) GPT-2 models have fewer parameters per layer so a GPT-2 layer ablation removes fewer parameters than a Pythia ablation.
4.2 Why are Language Models Robust to Layer-Wise Interventions?
We suspect that the robustness of language models can be partially attributed to the presence of residual connections [65] in the transformer architecture, which lead to increased redundancy [28, 45]. While skip connections were initially introduced to mitigate the vanishing gradient problem, investigations of deep residual networks (ResNets) found that residual connections facilitate "ensembling" within networks [34, 65]. We hypothesize that, as is the case for ResNets, residual connections in language models allow gradient descent to form ensembles of relatively shallow computational sub-networks. In doing so, networks avoid strong dependencies on individual paths, thereby increasing their resilience to layer-wise interventions. This hypothesis is supported by recent observations of self-repair mechanisms [44, 58] that demonstrate the existence of parallel computational paths.
5 Stages of Inference Hypothesis
We now discuss and provide tentative evidence for our hypothesis of four universal stages of inference in transformer language models. Two caveats apply to all of the following subsections: First, the boundaries between stages are fuzzy, and in practice, more than one stage can occur simultaneously. Second, these stages represent aggregate computational patterns, but the processing of any specific kind of token is likely to undergo more individualized dynamics (e.g., factual recall [46, 49]).
5.1 Stage 1: Detokenization
Given the extreme sensitivity of the model to first-layer ablations, we infer that the first layer is not so much a normal layer as it is an extension of the embedding. This is especially true for the Pythia model family due to the use of parallel attention, which implies that the first MLP layer is only a function of the current token. Consequently, by ablating the first layer, the rest of the network is blind to the immediate context and is thrown off distribution.
Immediately after computing this extended embedding, evidence from the literature suggests that the model concatenates nearby tokens that are part of the same underlying word [9, 20] or entity [49] (e.g., a first and last name). This operation integrates local context to transform raw token representations into coherent entities. In this way, the input is “detokenized” [15, 31].
Previous work has shown the existence of neurons that activate for specific -grams [31, 66]. Of course, to accomplish this, there must be attention heads that copy nearby previous tokens into the current token’s residual stream. More generally, if early layers are integrating local context, we would predict that early-layer attention heads pay disproportionately more attention to nearby tokens than later attention heads.
To test this, we compute the fraction of attention paid to tokens within the previous five positions of the present token. As can be seen in Figure 5, attention is indeed more local in the early layers. Moreover, the output norm of the attention ratio is higher than the output norm of the MLP layers in the first few layers [54].
5.2 Stage 2: Feature Engineering
Building upon the locally contextualized representations from stage 1, we hypothesize that a second region of the model performs "feature engineering" to construct features that could be useful for making downstream predictions, either for the next token or for future tokens in the context. While the kinds of features will vary by token type, there are many results in the literature that localize intermediate feature construction to the early to middle layers via patching [70] and probing [2] experiments.
For example, the model editing literature suggests that the MLPs in this region are important for factual recall [46, 25, 49]. This region is where probing accuracy for spatial and temporal features drastically but smoothly improves [30]. The features produced in this stage influence downstream predictions, as evidenced by steering and patching experiments on sentiment [62], truth [41], and zero-shot function execution [63] that peak in effectiveness in stage two layers. The representations and features formed in this stage are increasingly abstract, transitioning from shallower syntactic features to richer semantic features [15, 68, 38].
Logit Lens
However, in this stage, the features are simply produced, rather than consumed to make a concrete prediction. To show this, we perform a logit lens experiment [50, 12] where we apply the unembedding to the residual stream after every layer to estimate a model’s intermediate prediction. We then compute the entropy of this intermediate prediction and the KL divergence with the final prediction. As can be seen in Figures 10 and 6, there is very little progress towards making an actual prediction. For that, we require the next two stages.
5.3 Stage 3: Prediction Ensembling
After about the halfway point, the model must begin converting semantic features into concrete predictions for the next token. Given the robustness observed in Figure 3, we hypothesize that this is accomplished with a kind of ensembling. Ensembling in neural networks with residual connections is akin to having many subnetworks perform a "vote" for the output [65]. Therefore, interfering with a single member of an ensemble is unlikely to have a destructive effect.
Prediction Neurons
Previous work suggests that networks contain ensembles of “prediction" neurons, which act as probability promoters [66, 24, 32] and work in tandem with suppression neurons (Section 5.4). Following [32], we find prediction and suppression neurons by analyzing the output weights with the unembedding matrix . Prediction neurons exhibit a logit effect distribution with high kurtosis and positive skew, while suppression neurons show high kurtosis but negative skew. Here, is the output MLP weight for a given layer. Across 11 models, prediction neurons emerge around the midpoint, increasing in density towards the latter layers 6, before being outstripped by suppression neurons. To confirm their action, we study how much the model’s abstract representation changes as a function of prediction neuron density.
Change in Intermediate Prediction
To quantify how far the model’s representation is from the next prediction, we plot the "distance" (KL divergence) remaining for the model to reach its output distribution in Figure 6. We find that the rise in prediction neurons corresponds to a phase transition in the decrease of KL divergence. The number of prediction neurons peaks at around 85% through the model’s layers and decreases in the last few layers.
Phi-1 contains fewer prediction neurons than other Phi models and also has a lower slope in its KL divergence 6(c). GPT 6(a) and Phi 6(c) models exhibit more prediction neurons and steeper, smoother KL divergence slopes compared to Pythia 6(b). Surprisingly, Microsoft Phi models, which are known to outperform models with similar parameters, exhibit nearly 15% of prediction neurons per layer and 25% suppression neurons. This is 5-8x the density in GPT-2 and 3-7x the density in Pythia models, respectively. At around 90% through the model layers, however, the prediction neuron density decreases, while models continue approaching their final distribution, sometimes even accelerating 6(b). This suggests the action of other mechanisms, which we speculate is the final stage of inference.
5.4 Stage 4: Residual Sharpening
The subsiding of prediction neurons in the previous stage possibly suggests a new member of the ensemble, providing the final "push" to predicting the next token distribution. Our investigation reveals that the final layers of all models contain the highest density of suppression neurons, which may work to delete previously constructed features, suppress probabilities of invalid tokens, and/or calibrate the confidence in the final prediction.
Ensemble Bias
Prediction and suppression neurons both manipulate the residual stream and, as inverses of one another, can effectively perform each other’s functions. These neurons appear in different ratios and varying densities across the model. To study how these neurons sharpen the representation, we plot the logit lens entropy of the model. In certain models, such as Pythia (Figures 8(b) and 8(c)), the entropy sometimes increases in the final layers, suggesting overconfident predictions are blunted. In other words, the suppression neurons can either suppress tokens or features outside of the top-one to sharpen the distribution or suppress its confidence in the top token to flatten out the prediction distribution. This finding supports previous work, which suggests that models can shift away from the correct token to an incorrect token in the final layers [50, 61], and pruning or rank-reducing these layers can, in turn, improve performance [45, 28].
Final Layer
The intensity of suppression neurons, as seen in Figure 6, is localized in the final few layers of the model, where the quantity of suppression neurons outstrips predictive neurons. To quantify the intensity of this change, we measure the norm of the MLP output, where a larger norm suggests a greater contribution to the residual (Figure 7). Removal of the final layer or permuting its position results in the breakage of the model (Figure 3), analogous to the breakage observed in the first layers, during which the attention norm is the greatest (Figure 5). As a result, we speculate the importance of ordering in the first and last layers due to the magnitude of change they impart.
6 Case Studies
To integrate the stages of inference hypothesis with mechanistic descriptions of models, we present two case studies. First, we identify attention heads responsible for constructing multi-token words, known as subjoiner heads [20]. These heads help capture the context of a token for appropriate prediction, thus contributing to the detokenization and feature engineering stages of models. In the second case study, we provide evidence of the ensembling of prediction and suppression neurons. Through probing experiments, we demonstrate that multiple prediction and suppression neurons working jointly significantly outperform probes trained on individual neurons and sometimes even surpass the model’s performance.
6.1 Study 1: Attention on Four-Token Words
A crucial aspect of detokenization and feature engineering is building representations that integrate the context preceding a token. For a language model to understand and predict multi-token words, it must capture both the sentence context and the tokens that comprise a single word. To study these mechanisms, we construct a dataset with two classes: each consisting of 16 tokens, where in one class, the final 4 tokens form a word. We identify specific heads in the early layers of models that contribute solely to the construction of these multi-token words. As discussed in [20], these heads are called "subjoiner" heads. As illustrated in Figure 9, layer 2 head 5 of Pythia 2.8B moves information from earlier tokens to the final token of the word. The attention heads exhibit a consistent pattern, where attention decreases as tokens approach the final word. Specifically, the final token of the word attends most strongly to the first token, a feature absent in the baseline (shown on the left in Figure 9). This suggests at least one of many mechanisms by which models integrate local context, occurring at higher density in the first half of the models.
6.2 Study 2: Predicting the suffix -ing
Neurons performing prediction ensembling must work in tandem to predict the next token - akin to voting or operating in superposition. This suggests that multiple neurons working together may form a better prediction of the next token than a single neuron. To find evidence of this mechanism, we create a balanced dataset of two classes: tokens that do or do not end with the final token of "ing", all preceded by a context of 24 tokens. We train linear probes on the activation of 32 of the most active prediction and suppression neurons, both individually and in groups. We identify these neurons as outlined by [32], and provide examples of these neurons in GPT-2 XL (Figure 10(b)). See top the 36 neurons in the Appendix 16
Probing Results
We train two types of probes on activations at the penultimate token position of the dataset. First, we train 32 individual neuron probes and measure the classification accuracy (-ing/no -ing). We compare individual probes trained with the top-k neurons of the most accurate neurons, depicted by the line in Figure 10(a). We also note the mean model accuracy when predicting a token. Probes trained on suppression neurons, shown in yellow, resulted in the highest quality individual probes and performed similarly to the model itself, depicted by the dotted red line in Figure 10(a). Top-k probes trained with prediction neurons demonstrate even better accuracy than the average model prediction accuracy. Nonetheless, an individual neuron probe performs worse than any top-k probe, suggesting a critical role for ensembling in next-token prediction.
7 Concluding Remarks
Speculations of Duality
Our findings suggest that the second half of the model is, in some sense, dual to the first half. This was briefly suggested in [15] in the context of compound words being broken down token-wise in the early layer but rebuilt in later layers. Self-repair discusses erasure and anti-erasure pairs in the first half and second half of model [44, 58], as coupled attention heads in copy suppression [43]. Zooming out, the first half of models develop complex representation and the second half has the means to clear it 6. As seen by our experiments, interference with the first layer of models is also analogous to the interference of the final layer 3. While our study only suggests this we leave it for future work to investigate this further.
Model Mystery
A mystery occurs in the final layer of the GPT model, as suggested by Figure 5. Through our finds and suggested by [54], there is attention to MLP transition in all models, except an anomalous attention spike in the final layers of GPT, indicated by the . We speculate that this is caused by tied embedding and unembedding weights ( and ) during the training of GPT-2 models. By tying these weights, GPT might be going against the “duality" discussed above. Tied weights force the “input space" and “output space" to look identical. This tying might re-involve attention units as done in the early layers of all models; however, we leave this as a future avenue to explore.
Limitations
Our study does not conclusively identify the specific causes of differences between GPT and Pythia models, such as whether redundancy stems from dropout during training, structural variations in attention and MLP mechanisms, a greater number of layers, or a combination thereof. Furthermore, the study relies on aggregation over many tokens, which may average out effects that occur to specific token classes. Despite these uncertainties, our findings suggest the existence of phases of inference across these confounding factors.
Future Work
A natural place would be to address the aforementioned model limitations and other various references throughout the text. Further investigations into the stages of inference using Sparse Autoencoders (SAEs) may form a connection to provide more evidence for or against this hypothesis.
Contributions
VL conceived and led the study, performed all the analyses, and drafted the paper. WG and MT contributed to the experimental design and analytical methodology, and provided critical revisions of the paper, with WG additionally assisting in paper writing.
Acknowledgments
Eric Michaud, Josh Engels, Dowon Baek, Isaac Liao, and the rest of the lab for helpful feedback. Thank you to MIT Supercloud for providing the compute possible for the project [57].
Appendix A Cosine Similarity Analysis of Swapped Layers
Cosine Similarity Metrics
We collect three key metrics to compare a normal LLM to one with a set of adjacent layers swapped. First, self-similarity measures how much a layer retains its function after a swap, reflecting its "stubbornness." A high self-similarity score indicates that a layer continues to project similar contents to the residual stream, even after its position in the network has been changed. Second, index similarity assesses how closely the output of a swapped layer matches the output of the original layer it replaced. This metric serves as an indicator of a layer’s flexibility, with a high score suggesting that the layer can effectively assume the computational role of its predecessor, which could range from active processing to merely acting as a pass-through in the network. Lastly, adjacent similarity provides a baseline comparison by measuring the similarity in computations between adjacent layers in an unmanipulated model. This metric helps establish how similar or diverse the functions of neighboring layers are under normal conditions. By comparing these metrics across different stages of inference, we can gain insights into the commutativity of layers and the nature of the computations performed at each stage.
Cosine Similarity Results
Here we focus on Pythia 1.4B and GPT-2 XL, which contain a similar number of parameters (1.4B and 1.5B respectively). GPT-2 displays smoother trends compared to its Pythia counterpart while exhibiting similar overall patterns. We hypothesize that this is a result of differences in training dynamics (e.g., the use of dropout in GPT-2) and the fact that the GPT-2 model contains more layers. A larger number of layers presents greater opportunities for manipulating the output distribution and allows for more gradual changes. From an optimization perspective, this is analogous to taking smaller but more frequent gradient steps. Increasing the number of layers may also provide a means for greater redundancy in models, a key feature of GPT-style models that we discuss further below.
As seen in Figure 12, all model components maintain high degrees of self-similarity (denoted by the blue and red lines), suggesting that a component’s position does not significantly affect how it projects onto the residual stream when swapped. This finding has implications for how we interpret the remaining metrics. Another commonality across all plots is a significant change in metrics approximately halfway through the model, which we interpret as the separation between stages 2 and 3. Specifically, we observe a sharp decrease in index similarity and an increase in orthogonality between the swapped layer and its neighbors, suggesting a transition from iterative refinement to more specialized computations.
Attention Heads
Both models exhibit distinct patterns in attention-head behavior in the latter half of the network. In Pythia models, the attention head metrics converge to orthogonality, while in GPT-2 models, they converge to similarity. For Pythia, the self-similarity of attention heads decreases, indicating that they become less "stubborn" and more sensitive to their position in the network. In contrast, attention heads in GPT-2 models become increasingly redundant, with high self-similarity and index similarity scores. We hypothesize that this increased redundancy arises from the larger number of layers in GPT-2 models, which allows for a more gradual refinement of the output distribution. This finding has important implications for model design, suggesting that there may be an optimal number of layers given total parameters to balance computational efficiency and redundancy.
MLPs
The MLP components display two significant patterns across models. First, in the region corresponding to stage 2 of inference, we observe that the index similarity (teal line) is higher than both the adjacent similarity and the self-similarity scores. This pattern provides evidence for iterative inference, where a layer moved earlier in the computation has a projection onto the residual stream that overlaps more strongly with its previous neighbor than with its original position or its new neighbor. This overlap is more pronounced in Pythia models than in GPT-2 models, possibly because Pythia models have fewer layers to complete stage 2 of inference.
Second, in stage 3 of inference, both models demonstrate a convergence of all metrics except self-similarity toward orthogonality. The combination of high self-similarity (indicating stubbornness) and orthogonality to the replaced layer and the adjacent layers suggests a high degree of specialization in the MLPs of stage 3.
Appendix B Single Component Ablations Pythia 1.4B and GPT-2
Appendix C Top Prediction and Suppression Neurons
Appendix D Additional Empirical Details
All experimental code for future experiments is available at:
https://github.com/vdlad/Remarkable-Robustness-of-LLMs.
We make ubiquitous use of TransformerLens [48] to perform hooks and transformer manipulations.
For specificity, we utilize the following HuggingFace model names, and dataset. We do not change the parameters of the models from what they are described on the HuggingFace page.
| Name | HuggingFace Model Name |
| Pythia 410M | EleutherAI/pythia-410m-deduped |
| Pythia 1.4B | EleutherAI/pythia-1.4b-deduped |
| Pythia 2.8B | EleutherAI/pythia-2.8b-deduped |
| Pythia 6.9B | EleutherAI/pythia-6.9b-deduped |
| GPT-2 Small (124M) | gpt2 |
| GPT-2 Medium (355M) | gpt2-medium |
| GPT-2 Large (774M) | gpt2-large |
| GPT-2 XL (1.5B) | gpt2-xl |
| Phi 1 (1.3B) | microsoft/Phi-1 |
| Phi 1.5 (1.3B) | microsoft/Phi-1.5 |
| Phi 2 (2.7B) | microsoft/Phi-2 |
| The Pile | EleutherAI/the_pile_deduplicated |
All experiments described can be performed on a single NVIDIA A6000. We utilized 2 NVIDIA A6000 and 500 GB of RAM. To aggregate the metrics described in the paper, we run the model on 1 million tokens times, where is the number of layers. This takes on average 8 hours per model, per layer intervention (swapping and ablating). We save this aggregation for data analysis.
We utilize several conventional weight preprocessing techniques to streamline our calculations [48].
Layer Norm Preprocessing
Following [32], before each MLP calculation, a layer norm operation is applied to the residual stream. This normalizes the input before the MLP. The TransformerLens package simplifies this process by incorporating the layer norm into the weights and biases of the MLP, resulting in matrices and . In many layer norm implementations, trainable parameters and are included:
We "fold" the layer norm parameters into by treating the layer norm as a linear layer and then merging the subsequent layers:
Additionally, we then center reading weights. Thus, we adjust the weights as follows:
Centering Writing Weights
Because of the LayerNorm operation in every layer, we can align weighs with the all-one direction in the residual stream as they do not influence the model’s calculations. Therefore, we mean-center and by subtracting the column means of :